# 0.5pt (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v2

@@ 5.1 Experimental Setup
Dataset and metrics We test our method on the challenging 7 public scenes introduced with mip-NeRF 360 [3], comprising both bounded indoor and large unbounded outdoor settings. The data features a large amount of real images captured all around an object or scene. Given the high amount of views available, this dataset allows us to showcase the novel view quality in denser and sparser settings. For sparser setups, we select the training views by downsampling all images available (including the test ones) to preserve the coverage of the scene. We then select the test views among those test from the original set that were not selected as training images. Toward this end, we selected 30 training images (50 for stump and bicycle due to the convergence issues of the baselines when using 30 views). The exact data splits will be made available. We evaluate the novel views on the standard metrics and errors, i.e., PSNR, SSIM [53], and LPIPS [56].
sparser views denser views
Method PSNR SSIM LPIPS PSNR SSIM LPIPS
Instant-NGP [34] 20.76 0.659 0.301 25.87 0.812 0.162
+ Re-Nerfing [ours]* 21.40 0.682 0.270 25.88 0.814 0.161
PyNeRF [52] 22.65 0.746 0.182 27.41 0.874 0.094
+ Re-Nerfing [ours] 23.51 0.772 0.160 27.43 0.875 0.093
3DGS [26] 20.49 0.600 0.259 27.83 0.855 0.109
+ Re-Nerfing [ours]* 21.30 0.627 0.235 28.03 0.859 0.111
@@
Table 1: Rendering performance on the mip-NeRF 360 [3] dataset in denser (full) and sparser settings averaged over all scenes. The proposed Re-Nerfing is applied on the Instant-NGP [34], PyNeRF [52], and 3DGS [26] baselines, using the corresponding baseline to augment the views. * indicates that no uncertainty estimates were used.
@@
Prior Works and Baselines We showcase the effectiveness of our method by applying it on diverse novel view synthesis methods, namely PyNeRF [52], Instant-NGP [34], and 3DGS [26]. 3DGS uses an explicit representation, so there is no model uncertainty to estimate. Therefore, with 3DGS, we show only the impact of the simpler augmentation without masks.
@@
Implementation Details We train our method on an image resolution down-sampled by eight from the original image sizes for the available outdoor scenes and down-sampled by four for the indoor scenes, bringing both to a comparable image size. All models are trained using Nerfstudio [49]. For PyNeRF [52], we used the original implementation, which is compatible with Nerfstudio. Instead, for Gaussian Splatting [26] and Instant-NGP [34], we used their implementation available in Nerfstudio. We trained the models for 30k iterations with a batch size of 8192 rays for PyNerf and 1024 rays for Instant-NGP, and inherit all other losses and optimisation schemes of our baseline methods. The influence of the pseudo-views on the RGB loss is removed after 8k iterations for PyNerf and Gaussian-Splatting and after 200 iterations for Instant-NGP. We train all models on a single 24GB NVIDIA RTX 3090 or RTX 4090 GPU.
@@ 5.2 Quantitative Results
In Table 1, we report the results on the mip-NeRF 360 dataset [3] applying the proposed Re-Nerfing on Instant-NGP [34], PyNeRF [52], and 3D Gaussian Splatting [26], in both sparser and denser settings.
Re-Nerfing The proposed method brings notable improvements in sparser settings, increasing PSNR by 0.86 dB over PyNeRF and 0.64 dB over Instant-NGP, while improving SSIM and LPIPS as well. In sparser settings, the relatively low amount of views makes it difficult to disambiguate the scene geometry and properly render the novel views. In such settings, increasing the amount of images is particularly beneficial. Re-Nerfing generates such views with the baseline, thereby relying on the exact same original views. Yet, our iterative augmentation strategy brings valuable benefits, providing additional viewpoints on which the methods learn better representations of the scenes. For 3D Gaussian Splatting, our method brings a remarkable PSNR improvement of 0.81 dB over the baseline. Re-Nerfing aids in structuring the scene and reducing ambiguities. All sets of Gaussians were optimized from scratch with the same hyperparameters and settings. Despite their synthesized nature, adding more training views mitigates overfitting and acts as a regularizer. Such results are aligned with those of other iterative frameworks from other domains, where later models outperform earlier ones thanks to an implicit regularization effect occurring in the later stages [21, 10]. In denser settings, the benefit is reduced since the amount of original views is relatively high, with good coverage of the area of interest, allowing each method to reconstruct the scene significantly better. Therefore, in such setups, the synthesized views are rather close to the original ones, so the viewpoint advantage and the information gain are small. Nevertheless, Re-Nerfing slightly improves over each of its baselines, demonstrating its efficacy. Thanks to its simplicity, Re-Nerfing is widely applicable to both implicit and explicit methods, such as PyNeRF [52] and 3DGS [26], respectively.
@@
baseline, 2x iterations 19.68 0.633 0.274 17.93 0.653 0.331
baseline + 1 round of [ours] 21.08 0.686 0.241 18.09 0.676 0.299
baseline + 2 rounds of [ours] 21.89 0.720 0.214 18.36 0.693 0.285
baseline + 3 rounds of [ours] 22.47 0.748 0.191 18.41 0.699 0.281
baseline + 4 rounds of [ours] 22.63 0.756 0.184 18.47 0.704 0.275
baseline + 5 rounds of [ours] 22.76 0.761 0.178 18.64 0.708 0.272
Table 2: Rendering performance on the mip-NeRF 360 [3] counter scene in the sparser setting (30 views). The proposed Re-Nerfing is applied iteratively multiple times (rounds) and the effect of training the baseline for double the amount of iterations is also shown (2x). All other models in this work were trained for the same amount of iterations (30k).
Re-Nerfing as an iterative method In Table 2, we show the impact of our method when applying it iteratively. As described in Section 4.2, Re-Nerfing uses renderings from a first model $\mathbf{B}$ to augment the training data of a second model $\mathbf{D}$. To further improve the output quality, the process can be repeated iteratively, by using $\mathbf{D}$ to render the augmented views. We show this in Table 2, where at each iteration, the quality improves across the board for both pipelines, namely 3DGS [26] and PyNeRF [52]. This confirms the intuition that led to Re-Nerfing, for which novel view synthesis helps the task of novel view synthesis. While this varies from scene to scene, we notice that the performance saturates beyond 5 rounds. Overall, when applied to 3DGS [26], the proposed Re-Nerfing leads to a substantial boost of +3.45 PSNR, without using any extra data or external models, other than what is already available to the baseline.
Re-Nerfing vs. more training iterations Throughout this work, we optimized all models and Gaussians for 30k iterations to ensure fairness among the comparisons. By being multi-stage, applying the proposed Re-Nerfing, increases the total training time, e.g., by 2x considering a single round of Re-Nerfing, albeit resetting the optimization every time. Therefore, in Table 2, we assess the impact of doubling the number of iterations for the baselines. This results in a total number of iterations and training time equivalent to 1 round of Re-Nerfing. While doubling the iteration count improves the output quality of the baseline, this remains significantly below applying Re-Nerfing for 1 round. Such comparison showcases the significant benefits of the proposed method, which introduces synthetic novel views among the original ones.
@@
PyNeRF [52] 22.42 0.632 0.218
@@
Table 3: View selection strategies on the sparser setting of the stump scene of the mip-NeRF 360 [3] dataset (50 views). All models are based on PyNeRF [52]. Augmentation factors $N$ from 2 to 11 are evaluated. 11x corresponds to the $\alpha$ factors .05-.1-.2-.3-.4-.5-.6-.7-.8-.9-.95.
@@
Trade-offs with pseudo-views poses In Table 3, we evaluate different poses for the novel view selection on PyNeRF [52]. All interpolate across pairs of the original views, but each does it at different locations via the interpolation factors $\alpha_{j}$ (e.g., 0.33 and 0.75). Remarkably, all augmentations are beneficial over PyNeRF, bringing a gain in PSNR between +0.9 and +1.3 dB. The results show the trade-offs depicted in Figure 3: the furthest from the original views (e.g., .33-.66 is further than 0.20-0.80), the lower the quality of the novel views, and the smaller the performance improvement. However, the opposite trade-off occurs, too: the closer from the initial images (e.g., 0.10-0.90 are closer than 0.20-0.80), the less amount of new information to gain for the multi-view optimization. Considering these trade-offs, it is better to sample more pseudo-views closer to the original ones while still including further viewpoints to add valuable information. We do so with the 7x configuration, where the increments along the interpolation are smaller around the original views and increase further away from 0.1 to 0.2.
@@
Impact of the augmentation factor In Table 3, we also explore different augmentation factors $N$ (e.g., 2x and 5x). The results show that adding more images (i.e., higher factors) leads to improvements despite the introduction of sub-optimal, synthesized views. Doubling the views (2x) brings the largest relative gain with +0.9 dB PSNR. Then, the performance further improves when the augmentation factor is increased. However, at high $N$ factors (e.g., 11x), benefits reduce as the influence of the many synthetic views overcomes the few original images. 7x provides a good balance and leads to the best scores.
Ablation study In Table 4, we report an ablation study for both PyNeRF and Instant-NGP in the sparser setting (30 views). Training again with the synthetic views produced by the baseline (with $N=7$) degrades the performance for all metrics for PyNeRF, while it improves for Instant-NGP. The degradation for PyNeRF is due to the lower quality signal of 6x more views than the reliable signal of the original views. Specifically, the model learns the artifacts of the augmented views and underperforms. Instead, for Instant-NGP, the quality of the augmented views was higher, leading to improvements. Nevertheless, removing the images from the sampling pool once the scene structure has been mainly learned (stop) prevents both models from overfitting to the artifacts and leads to better results on all metrics, with large improvements for PyNeRF. Furthermore, masking out the uncertain regions of such synthetic views improves even further across the board for both methods. This shows once again how synthetic views of the same scene are beneficial aid the novel view synthesis task. In particular, the results confirm the importance of increasing the scene coverage and that of having a good training signal, testifying the need for the uncertainty-based masks and our augmentation strategy.

### v3

@@ 4.1 Experimental Setup
Dataset and metrics We test our method on the challenging 7 public scenes introduced with mip-NeRF 360 [3], comprising both bounded indoor and large unbounded outdoor settings. Additionally, we experimented with the 8 scenes of the LLFF dataset [36], which includes smaller, bounded scenes. The former features a large amount of real images captured all around an object or scene, while the latter has fewer views of forward-facing scenes. So, the two assess complementary aspects. For sparsifying the mip-NeRF 360 dataset, we select the training views by sampling from all available images (including the test ones) while preserving as much coverage of the scene as possible. We then select the test views among those test views from the original set that were not selected as training images. Toward this end, we selected 30 training images (50 for stump and bicycle due to convergence issues of the baselines when using 30 views). The exact data splits will be made available. With LLFF, we experimented with the highly sparse setting with just 3 views available, following the evaluation paradigm of RegNeRF [40]. We evaluate the test views on the standard metrics and errors, i.e., PSNR, SSIM [60], and LPIPS [63].
Method PSNR SSIM LPIPS
Instant-NGP [38] 20.76 0.659 0.301
+ Re-Nerfing [ours] 21.35 0.680 0.273
PyNeRF [59] 22.65 0.746 0.182
+ Re-Nerfing [ours] 23.51 0.772 0.160
3DGS [28] 20.49 0.600 0.259
+ Re-Nerfing [ours] 21.30 0.627 0.235
@@
Table 1: Rendering performance on the mip-NeRF 360 [3] dataset in sparse view settings averaged over the scenes. The proposed Re-Nerfing is applied on the Instant-NGP [38], PyNeRF [59], and 3DGS [28] baselines, using the corresponding baseline to augment the views. Uncertainty estimates are unavailable for 3DGS.
@@
Prior Works and Baselines We showcase the effectiveness of our method by applying it on diverse NVS methods, namely PyNeRF [59], Instant-NGP [38], RegNeRF [40], and 3DGS [28]. 3DGS uses an explicit representation, so there is no model uncertainty to estimate. RegNeRF, on the other hand, is implemented in JAX [6], for which no implementation of Bayes’Rays [25] exists. Therefore, with 3DGS and RegNeRF, we show only the impact of our augmentations without uncertainty masks.
@@
Implementation Details We train our method on an image resolution down-sampled by eight from the original image sizes for the available outdoor scenes and down-sampled by four for the indoor scenes of the mip-NeRF360 dataset, bringing both to a comparable image size. For the LLFF dataset we down-sample all images by eight, following RegNeRF [40]. All models are trained using Nerfstudio [56], except for RegNeRF where we use the original implementation in JAX [6]. For PyNeRF  [59], Gaussian Splatting [28] and Instant-NGP [38], we used their implementation available in Nerfstudio. We trained all Nerfstudio models for 30k iterations from scratch with a batch size of 8192 rays for PyNerf and 1024 rays for Instant-NGP and inherited all other losses and hyperparameters of the baseline methods. For RegNeRF we followed their training paradigm and hyperparameters. The influence of the synthesized views on the optimization is removed after 8k iterations for PyNerf and Gaussian-Splatting, after 5K iterations for RegNeRF and after 200 iterations for Instant-NGP. We train all models on a 24GB NVIDIA RTX 3090 or 4090 GPU.
@@ 4.2 Quantitative Results
In Table 1, we report the results on the mip-NeRF 360 dataset [3] applying the proposed Re-Nerfing on Instant-NGP [38], PyNeRF [59], and 3D Gaussian Splatting [28] in sparse view settings. In Table 2, we report the results on the LLFF dataset [36] applying Re-Nerfing on RegNeRF [40] in highly-sparse 3 view settings.
3 Views PSNR SSIM LPIPS
baseline: mip-NeRF [2] 14.62 0.351 0.495
+ RegNeRF [40] 19.08 0.685 0.148
+ Re-Nerfing, 1 round [ours] 19.35 0.713 0.133
+ extra round, 2nd [ours] 19.47 0.724 0.126
+ extra round, 3rd [ours] 19.52 0.725 0.127
Table 2: Highly sparse settings with only 3 views on the LLFF dataset [36]. We apply ours on top of RegNeRF [40]. Then, we iteratively apply ours via the extra rounds. The results are averaged over all LLFF scenes.
Re-Nerfing as flexible add-on The proposed method brings notable improvements across various settings, increasing PSNR by 0.86 dB over PyNeRF, 0.81 dB over 3DGS, 0.64 dB over Instant-NGP, and 0.27 dB over RegNeRF already with a single iteration, while improving SSIM and LPIPS as well. Despite their synthesized nature, adding more training views mitigates overfitting and acts as a regularizer. In Table 2, in particular, we combine Re-Nerfing with the regularization of RegNeRF, showing how Re-Nerfing can adapt to any NVS method. This is because it is an add-on and not an alternative to them. Overall, the efficacy of our method on both implicit (NeRF-based) and explicit (3DGS) scene representations is particularly notable as it is achieved without using extra data, additional models, or supervision, but only with already available information.
Re-Nerfing as an iterative method In Tables 2 and 3, we show the impact of our method when applying it iteratively. As described in Section 3.3, Re-Nerfing uses renderings from a first model $\mathbf{B}$ to augment the training data of a second model $\mathbf{D}$. To further improve the output quality, the process can be repeated iteratively, by using $\mathbf{D}$ to render the augmented views. We show this in the tables, where, at each iteration, the quality improves across the board for the various pipelines and settings. This confirms that NVS helps the task of NVS. While the improvement varies from scene to scene, we notice that the performance saturates beyond 3-5 rounds. Overall, the proposed Re-Nerfing leads to a substantial boost of +3.45 PSNR over 3DGS [28], without using any extra data or external models, other than what is already available to the standard 3DGS.
@@ 3 Views PSNR SSIM LPIPS
+ [ours]: 1 21.08 0.686 0.241 18.09 0.676 0.299
+ [ours]: 2 21.89 0.720 0.214 18.36 0.693 0.285
+ [ours]: 3 22.47 0.748 0.191 18.41 0.699 0.281
+ [ours]: 4 22.63 0.756 0.184 18.47 0.704 0.275
+ [ours]: 5 22.76 0.761 0.178 18.64 0.708 0.272
Table 3: Rendering performance on the mip-NeRF 360 [3] counter scene in the 30-view setting. Re-Nerfing is applied iteratively across multiple rounds, from 1 to 5.
@@
1x: PyNeRF [59], baseline 22.42 0.632 0.218
1.08x: 4 random syn.views 22.58 0.653 0.205
1.08x: 4 test syn.views 22.64 0.663 0.196
@@
Table 4: View selection strategies on the stump 50-view setting of the mip-NeRF 360 [3] dataset. Models based on PyNeRF [59]. Our augmentation factors $N$ from 2 to 11 are evaluated.
Why it works As described in Section 3.1, NeRF’s and 3DGS’ optimization is similar to SfM approaches, as matching pixel features are triangulated in 3D. Especially in sparse settings, this is highly under-constrained with many degenerate solutions. This can lead to local minima, which fit well with the training views but not the test views due to the inconsistent geometry. With Re-Nerfing, we build upon this (potentially sub-optimal) initial optimization and render synthetic views as extra constraints for later rounds. It is widely known that more viewpoints help NVS. This same concept drives Re-Nerfing, too, albeit with synthesized views. Thanks to these extra constraints, some degenerate solutions are removed and the optimization converges to a better 3D structure. This can be seen in the supplementary video, with smoother rendered depth maps. Through an ablation study on 3DGS [28], Table 5 shows the importance of: early optimization stages for the scene geometry and the impact of the pseudo views (syn.views) on that, removing the pseudo views as the scene geometry converges to not learn wrong details from them, and resetting the training at each of the Re-Nerfing rounds. Specifically, adding the synthetic views without resetting the optimization (3rd row) does not bring any benefits and worsens LPIPS. Resetting and keeping the rendered views until the end (4th row) improves, but removing them leads to better details (last row, ours). On the contrary, adding the views after the scene geometry has already been captured (5th row) leads to mixed results compared to the baselines. While Re-Nerfing increases the training iterations, e.g., by 2x in the table, simply doubling the iterations performs worse (2nd row). All together, Re-Nerfing leads to a better global solution without adding extra info over the baseline. This is similar to other iterative frameworks from other domains [23, 10], where resetting the optimization at each round leads to better solutions.
sparse views PSNR SSIM LPIPS
baseline: 3DGS [28], 30k iter. 19.31 0.609 0.296
2x longer training (60k iter.) 19.68 0.633 0.274
no reset, add syn.views at 30k 19.67 0.631 0.333
reset, add and keep syn.views 20.15 0.654 0.291
reset, add syn.views at 15k 19.79 0.637 0.325
reset, remove syn.views at 8k 21.08 0.686 0.241
Table 5: Ablation study. 30-view setting, counter scene of the mip-NeRF 360 dataset [3] with 3DGS [28]. The last row represents the proposed method.
Figure 3: Qualitative results on cropped images from the test set of the mip-NeRF 360 and LLFF datasets. 3DGS [28] and Instant-NGP [38] were trained on the mip-NeRF 360 dataset in a 30 view setting, while RegNeRF [40] was trained on LLFF with only 3 views. Qualitative results of PyNeRF [59] can be found in the supplementary material.
@@
Trade-offs with pseudo-views poses In Table 4, we evaluate different poses for the novel view selection on PyNeRF [59], exploring different interpolations across pairs of the original views. Remarkably, all augmentations are beneficial, improving PSNR by 0.9-1.3 dB. The results show the trade-offs further described in the supplementary material: the furthest from the original views (e.g., .33-.66 is further than 0.20-0.80), the lower the quality of the novel views, and the smaller the performance improvement. However, the opposite trade-off occurs, too: the closer to the initial images, the less amount of new information to gain for the multi-view optimization. Therefore, it is better to sample more pseudo-views closer to the original ones while still including further viewpoints for extra information, as for 7x.
@@
Impact of the augmentation factor In Table 4, we also explore different augmentation factors $N$. Higher factors are more beneficial despite the introduction of more sub-optimal, synthesized views. Doubling the views (2x) brings the largest relative gain with +0.9 dB PSNR. The performance further improves as the factor is increased. However, at high $N$ factors (e.g., 11x), benefits reduce as the influence of the many synthetic views overcomes the few original images, with 7x striking a good balance.
Optimizing for specific views In Table 4, we show how we can optimize for specific views, by augmenting with 4 test views synthesized by the baseline. We compare this with augmenting with 4 randomly interpolated views and evaluate on the viewpoint of the 4 test views. While adding random views helps, specifically synthesizing the viewpoints of interest (e.g., the test views) shows higher margins on all metrics. In practice, Re-Nerfing allows to improve on viewpoints of interest, which can be beneficial when the available views lead to a sub-optimal quality in specific areas. In such settings, Re-Nerfing can densify the views around such areas and ultimately improve the outputs.
4.3 Qualitative Results
Figure 3 shows examples of synthesized images with 3DGS [28], RegNeRF [40], and Instant-NGP [38], with and without our method. As already seen in the quantitative results (Section 4.2), the output quality of the proposed Re-Nerfing improves significantly upon the baselines, especially in terms of the reduction of floating artifacts. In particular, the addition of our method sharpens the scene and better preserves the small details as can be seen for RegNeRF in Figure 3. Remarkably, the proposed method is able to remove dominant artifacts in the renderings by Instant-NGP and 3DGS, despite using the same original views, architecture, and hyperparameters as the baselines. The augmented views help disambiguating the scene geometry, allowing all models to deliver better results by converging to a better solution.
The supplementary material includes additional details and results, such as the performance on each scene and extra qualitative examples in a video. There we also show how the proposed Re-Nerfing can densify the SfM point cloud that can be used to initialize 3DGS.
Limitations Our method relies on the first stage to provide reasonable renderings and a decent estimate of the scene geometry, so it is challenging to enhance poorly rendered scenes, e.g., the stump scene with Instant-NGP [38], which could not converge. The proposed method also relies on the uncertainty estimates to discard artifacts and low-quality areas. Therefore, better uncertainty estimates would be beneficial. Uncertainty calibration of the threshold $\mu$ could help refine the estimates while preserving the signal in higher-quality areas.
5 Conclusion
This work enhances NeRFs and 3DGSs by leveraging their own novel view synthesis capabilities. The introduced Re-Nerfing is a general approach that synthesizes novel views to improve the scene coverage and augment the training data for a new model. To provide a valuable training signal for the second model, Re-Nerfing computes the NeRF’s uncertainty for each augmented view to discard the uncertain regions and reduce artifacts. As shown with various pipelines, our method yields improvements in both reconstruction and rendering quality in sparse scenes, mitigating the need for extensive captures of the scene.
[…]
Supplementary Material
Appendix A Appendix
In this supplementary material, we complement the main paper with additional details and results. Specifically, we illustrate and discuss the trade-offs due to the synthetic view poses (Section A.1), we experiment with and discuss the impact of our method on the SfM point cloud (Section A.2), we present the result in dense-view settings (Section A.3), we discuss an extra ablation study (Section A.4), we show more qualitative results, also through a video (Section A.5), and we present the quantitative results by scene (Section A.6).
Method PSNR SSIM LPIPS
Instant-NGP [38] 25.87 0.812 0.162
+ Re-Nerfing [ours] 25.88 0.814 0.161
PyNeRF [59] 27.41 0.874 0.094
+ Re-Nerfing [ours] 27.43 0.875 0.093
3DGS [28] 27.83 0.855 0.109
+ Re-Nerfing [ours] 28.03 0.859 0.111
Table 6: Rendering performance on the mip-NeRF 360 [3] dataset in dense view settings averaged over the scenes. The proposed Re-Nerfing is applied on the Instant-NGP [38], PyNeRF [59], and 3DGS [28] baselines, using the corresponding baseline to augment the views. Uncertainty estimates are unavailable for 3DGS.
A.1 More on the Views Positioning Trade-Offs
In Figure 5, we illustrate the trade-offs associated with the relative positioning between the original views and the synthesized ones. Considering two existing views, the quality of the views synthesized between them varies as follows: the closer the synthetic views are to the original views, the higher their quality because the viewpoint change is small. Instead, the further from the original views (e.g., towards 0.5), the worse the quality becomes (e.g., more artifacts) as the viewpoint shift is more significant and the uncertainty increases. However, synthesizing views only in the vicinity of the original ones is not particularly beneficial as the views are too similar to the original ones and, as such, do not help disambiguate the scene geometry. Therefore, the views should also be synthesized at further locations. With Re-Nerfing, to mitigate the quality degradation, we remove the artifacts within the synthesized views (i.e., occurring more at the further poses) by filtering out the uncertain regions. This allows us to synthesize many views throughout the interpolated poses. We sample more poses closer to the original views (i.e., 0.0 and 1.0) and fewer poses in the middle. Specifically, we synthesize in the 7x configuration at the interpolation factors 0.1, 0.2, 0.4, 0.6, 0.8, and 0.9. This trade-off is explored in Table 4.
Figure 4: Qualitative results on cropped images from the test set of the mip-NeRF 360 dataset [3] showcasing PyNeRF [59] in sparse-view settings, with and without the proposed Re-Nerfing.
3DGS [28] (SfM, Hloc+Superpoint) Re-Nerfing (SfM, Hloc+Superpoint)
Scene - sparse PSNR SSIM LPIPS # Points PSNR SSIM LPIPS # Points
bonsai 26.19 0.847 0.131 6683 26.52 0.864 0.126 20274
counter 23.29 0.766 0.167 8032 23.32 0.773 0.165 30855
kitchen 24.27 0.849 0.098 3914 24.22 0.852 0.099 13716
room 24.82 0.829 0.162 4304 24.73 0.833 0.167 16216
garden 24.33 0.773 0.089 4873 24.87 0.793 0.081 13646
stump 21.97 0.528 0.230 7337 22.36 0.568 0.228 38891
bicycle 20.04 0.482 0.329 18831 20.38 0.514 0.301 114554
average 23.56 0.725 0.172 7711 23.77 0.742 0.167 35450
Table 7: Impact of Re-Nerfing on the SfM point cloud and 3DGS [28] initialized with it. SfM is run for all with HLoc [50] with Superpoint [16] features. For Re-Nerfing, SfM is run with the synthesized views, too.
A.2 Impact on the SfM Point Cloud
In Table 7, we investigate the impact of Re-Nerfing on the SfM point cloud. For these experiments, we used 3DGS [28] initialized with the SfM point cloud constructed by the SfM pipeline of Nerfstudio [56] using Hloc [50] and Superpoint [16] from the sparse settings. For the second stage of Re-Nerfing, we followed the same protocol of Section 3 but ran the aforementioned SfM-pipeline on the synthetic and real images combined. Remarkably, despite synthetic views, not only did Re-Nerfing improve over the standard 3DGS but also densified the point cloud by a substantial 4.6x on average. This significant difference in the number of points (# points) is evidence of the effectiveness of NVS and Re-Nerfing to aid and simplify the triangulation of the scene geometry, which is the foundation of our work. In particular, the bicycle scene has more points than the other ones because of its higher image resolution. Thanks to a much denser point cloud as initialization and the added synthesized views, the proposed Re-Nerfing reaches a better solution, as shown by the metrics.
A.3 Dense-View Settings
Throughout this work, we experimented with challenging sparse-view settings. There, the limited views available prevent a good triangulation of the scene geometry. In Table 6, we show the impact of our Re-Nerfing on Instant-NGP [38], PyNeRF [59], and 3DGS [28] in dense-view settings. Therefore, in these experiments, we use all the available training views. Remarkably, despite the already many views available, by adding synthesized ones, the proposed Re-Nerfing does not degrade the performance and even slightly improves it. Nearly equal results are seen for Instant-NGP and PyNeRF, while a +0.2db PSNR is brought over 3DGS. This shows the flexibility of Re-Nerfing.